OpenStudy (thadyoung):

describe how the graphes of y=/x/ and y=/x/-15 are related A:the graphs have the same shape.The y-intersept of y=/x/ is 0, and the x-intercept of the second graph is -15 B: the graphs have the same y-intercept. the second graph is steeper then y=/x/ C:the two graphs are the same D:the graphes have the same shape.the y-intercept of y=/x/is 0, and the y-intercept of the second graph is -15

OpenStudy (mathmale):

Hi, Thad, Have you any experience in graphing y = |x|, the absolute value function? Can you graph y=x^2 quickly and confidently? If so, how would the graph of y=x^2-3 appear different from that of y=x^2? Your responses will help me know better where to start in helping you with this problem.

OpenStudy (thadyoung):

i dont know where to start at all

OpenStudy (thadyoung):

history and bio are my things maybe english but math? nope lol my brain freezes up and i cant think

OpenStudy (mathmale):

If history, bio and English are more your thing, at least that says you're likely a pretty good reader. Thad, have you looked up "absolute value function" in your textbook? There might be an alternative definition of that function, as well as a graph of it.

OpenStudy (thadyoung):

im in online schooling and i might be able to go back a few assignments and re read some lessons do you think that this would be the best thing to do?

OpenStudy (mathmale):

Only if you have a definite focus in doing so. Let me ask again: Are you at all familiar with the graph of the quadratic function y=x^2 (whose shape is that of a parabola)? And a new question: have you heard of "horizontal translation" or "horizontal shifting" of graphs, or of "vertical shifting?"

OpenStudy (thadyoung):

i have heard of the shifting and translations yes

OpenStudy (mathmale):

OK. I'm going to assume that you'll recognize the graph of the parabola y=x^2. I'll sketch that graph for you. Then I'll shift that graph vertically DOWN by 2 units, as an example of vertical translation. hang in there.

OpenStudy (mathmale):